# MULTIGRADED BETTI NUMBERS OF VERONESE EMBEDDINGS

### v1

@@ 1 INTRODUCTION
A central open question in the study of syzygies is to determine the Betti table of $\mathbb{P}^{m}$ under the $d$-uple Veronese embedding. Veronese syzygies have been the focus of a great deal of attention, e.g. [zbMATH07330979, zbMATH07124911, zbMATH06609392, zbMATH06117694, zbMATH06620296, arXiv:2311.03625, zbMATH01587223, zbMATH05268889]. The case $m=1$ is well understood, but even the case $m=2$ is wide open. For the $m=2$ case, in a major computational effort, [zbMATH07330979] recently computed all multigraded Betti numbers up to the sixth Veronese (and some up to the eighth). Our goal in this paper is to better understand the multigraded Betti numbers of Veronese embeddings of projective spaces. In particular, we prove new vanishing and non-vanishing results for these multigraded Betti numbers. Related vanishing theorems were proved with other methods by Castryck, Lemmens, and Hering [CastryckLemmensHering]. Their bounds are incomparable to ours.
What are the multigraded Betti numbers of Veronese embeddings of projective spaces? Fix integers $d\geq 2$ and $m\geq 2$. We consider the set $\mathcal{A}=\{\mathbf{a}\in\mathbb{N}^{m+1}\mid|\mathbf{a}|=d\}$, where $|\mathbf{a}|\mathrel{\mathop{:}}=\mathbf{a}_{0}+\mathbf{a}_{1}+\cdots+\mathbf{a}_{m}$ denotes the sum of all coordinates of $\mathbf{a}$. Let $\mathbf{a}^{i}$ be the $i$-th element of $\mathcal{A}$ with respect to the lexicographical order, let $n\mathrel{\mathop{:}}=\binom{d+m}{m}$ be the cardinality of $\mathcal{A}$, and denote $\mathbb{N}\mathcal{A}\subset\mathbb{N}^{m+1}$ be the semigroup generated by $\mathcal{A}$.
@@
Let $S=\mathbb{C}[x_{1},\dots,x_{n}]$. The homomorphism $\pi$ defines a multigrading on $S$, where $\deg(x_{i})=\mathbf{a}^{i}$. And this induces a $\mathbb{C}$-algebra homomorphism $\hat{\pi}\colon S\to\mathbb{C}[\mathbb{N}\mathcal{A}]$. Note that $\mathbb{C}[\mathbb{N}\mathcal{A}]$ is the $d^{\text{th}}$ Veronese subring of $\mathbb{C}[y_{0},y_{1},\dots,y_{m}]$, and it has an $S$-module structure via the map $\hat{\pi}$. The $S$-module $\mathbb{C}[\mathbb{N}\mathcal{A}]$ is multigraded, with graded components given by
@@
Our goal is to study the multigraded Betti numbers $\beta_{p,\mathbf{b}}=\beta_{p,\mathbf{b}}(\mathbb{C}[\mathbb{N}\mathcal{A}])$ of the multigraded $S$-module $\mathbb{C}[\mathbb{N}\mathcal{A}]$. Geometrically, this corresponds to studying the multigraded Betti numbers of the coordinate ring of $\mathbb{P}^{m}$ under the $d$-uple Veronese embedding $\mathbb{P}^{m}\to\mathbb{P}^{n-1}$.
@@
has the upper bound $A_{j}-1$. It has the lower bound $\tilde{l}_{j}+1$ if we assume $j\geq d+1$. Here, $\operatorname{pr}_{s}$ denotes the projection onto the $s+1$-th coordinate.
We show that the upper bounds are optimal. In particular, for each $1\leq p\leq\binom{d+m-1}{m}+m-2$, there is a $\mathbf{b}\in\mathbb{N}^{m+1}$ with $|\mathbf{b}|=d(p+1)$ and $\mathbf{b}_{0}=A_{p+1}-1$ such that $\beta_{p,\mathbf{b}}\neq 0$.
Applying Forman’s discrete Morse theory [zbMATH02041483], we compute the exact Betti numbers when $\mathbf{b}_{0}=A_{p+1}-1$ and state the result as follows.
@@ Acknowledgements
We thank Milena Hering and the authors of [zbMATH07330979] for inlightening conversations and encouragement.
@@ 2.1 Hochster’s formula
In this subsection, we provide background on the simplicial complex $\Delta_{\mathbf{b}}$ and Hochster’s formula. The references for these are [zbMATH02190625, Section 9.1] and [zbMATH00835749, Section 12.C].
@@
$\beta_{p,\mathbf{b}}=\dim_{\mathbb{C}}\tilde{H}_{p-1}(\Delta_{\mathbf{b}};\mathbb{C})$.
@@ 2.2 Discrete Morse theory
In this subsection, we provide background on discrete Morse theory (cf. [zbMATH02041483, zbMATH07258477]).
@@
A discrete Morse function on $\Delta$ [zbMATH02041483, Definition 2.1] is a function which, roughly speaking, assigns higher numbers to higher-dimensional simplices, with at most one exception, locally, at each simplex. Once a discrete Morse function $f$ is given, the discrete gradient vector field of $f$ is defined as a collection $W$ of pairs $\{\alpha^{(q)}<\beta^{(q+1)}\}$ of simplices, where $\{\alpha^{(q)}<\beta^{(q+1)}\}$ is in $W$ if and only if $f(\beta)\leq f(\alpha)$. A simplex is critical [zbMATH02041483, Definition 2.3] if and only if it is not in $W$.
Theorem 2.5 ([zbMATH02041483, Theorem 3.5]).
@@ 6 GENERALIZATION TO HIGHER DIMENSIONAL PROJECTIVE SPACE
$D\mathrel{\mathop{:}}=\Big\{I\in\binom{\{\binom{r+m-2}{m}+1,\dots,\binom{r+m-1}{m}\}}{p+1-\binom{r+m-2}{m}}\mid\mathbf{b}_{t}-\displaystyle\sum_{i=1}^{\binom{r+m-2}{m}}\mathbf{a}^{i}_{t}-1\leq\displaystyle\sum_{i\in I}\mathbf{a}^{i}_{t}\leq\mathbf{b}_{t}-\displaystyle\sum_{i=1}^{\binom{r+m-2}{m}}\mathbf{a}^{i}_{t}\quad\forall 1\leq t\leq m-1\Big\}$
@@ 7 Optimality of the Bounds $A_{p+1}$
The articles [zbMATH06117694, zbMATH06609392] show that $\beta_{p,1}\neq 0$ for $1\leq p\leq\binom{d+m-1}{m}+m-2$. In the following, we give a combinatorial proof and show that, on the row $1$ of the Betti table, the bounds $A_{p+1}$ for the multigraded Betti numbers are sharp.
@@
Christian Haase: Mathematik, Freie Universität Berlin; haase@math.fu-berlin.de
Zongpu Zhang: Mathematik, Freie Universität Berlin; zongpuzhang@zedat.fu-berlin.de

### v2

@@ 1 INTRODUCTION
A central open question in the study of syzygies is to determine the Betti table of $\mathbb{P}^{m}$ under the $d$-uple Veronese embedding. Veronese syzygies have been the focus of a great deal of attention, e.g. [BEG+20, CCD+19a, EEL16, EL12, GM16, KEM23, OP01, RUB04]. The case $m=1$ is well understood, but even the case $m=2$ is wide open. For the $m=2$ case, we do know which coarsely graded Betti numbers vanish [APR02, GRE84, BIR95]:
$\beta_{p,0}\neq 0$ $\iff p=0$
$\beta_{p,1}\neq 0$ $\iff 1\leq p\leq\binom{d+1}{2}$ (1)
$\beta_{p,2}\neq 0$ $\iff 3d-2\leq p\leq\binom{d+2}{2}-3\,.$
Here, we focus on the multigraded Betti numbers. In a major computational effort, [BEG+20] recently computed all multigraded Betti numbers up to the sixth Veronese (and some up to the eighth). Our goal in this paper is to better understand the multigraded Betti numbers of Veronese embeddings of projective spaces. In particular, we prove new vanishing and non-vanishing results for these multigraded Betti numbers. Related vanishing theorems were proved with other methods by Castryck, Lemmens, and Hering [CLH19b]. Their bounds are incomparable to ours.
What are the multigraded Betti numbers of Veronese embeddings of projective spaces? Let $k$ be any field. Fix integers $d\geq 2$ and $m\geq 2$. We consider the set $\mathcal{A}=\{\mathbf{a}\in\mathbb{N}^{m+1}\mid|\mathbf{a}|=d\}$, where $|\mathbf{a}|\mathrel{\mathop{:}}=\mathbf{a}_{0}+\mathbf{a}_{1}+\cdots+\mathbf{a}_{m}$ denotes the sum of all coordinates of $\mathbf{a}$. Let $\mathbf{a}^{i}$ be the $i$-th element of $\mathcal{A}$ with respect to the lexicographical order, let $n\mathrel{\mathop{:}}=\binom{d+m}{m}$ be the cardinality of $\mathcal{A}$, and denote $\mathbb{N}\mathcal{A}\subset\mathbb{N}^{m+1}$ be the semigroup generated by $\mathcal{A}$.
@@
Let $S=k[x_{1},\dots,x_{n}]$. The homomorphism $\pi$ defines a multigrading on $S$, where $\deg(x_{i})=\mathbf{a}^{i}$. And this induces a $k$-algebra homomorphism $\hat{\pi}\colon S\to k[\mathbb{N}\mathcal{A}]$. Note that $k[\mathbb{N}\mathcal{A}]$ is the $d^{\text{th}}$ Veronese subring of $k[y_{0},y_{1},\dots,y_{m}]$, and it has an $S$-module structure via the map $\hat{\pi}$. The $S$-module $k[\mathbb{N}\mathcal{A}]$ is multigraded, with graded components given by
@@
Our goal is to study the multigraded Betti numbers $\beta_{p,\mathbf{b}}=\beta_{p,\mathbf{b}}(k[\mathbb{N}\mathcal{A}])$ of the multigraded $S$-module $k[\mathbb{N}\mathcal{A}]$. Geometrically, this corresponds to studying the multigraded Betti numbers of the coordinate ring of $\mathbb{P}^{m}$ under the $d$-uple Veronese embedding $\mathbb{P}^{m}\to\mathbb{P}^{n-1}$.
@@
has the upper bound $A_{j}-1$. We show that the upper bounds are optimal (cf. Theorem 7.1). It has the lower bound $\tilde{l}_{j}+1$ if we assume $j\geq d+1$. Here, $\operatorname{pr}_{s}$ denotes the projection onto the $s+1$-th coordinate.
Applying Forman’s discrete Morse theory [FOR02], we compute the exact Betti numbers when $\mathbf{b}_{0}=A_{p+1}-1$ and state the result as follows.
@@
The simplicial complex $\Delta_{\mathbf{b}}$ has been studied under the name bounded degree (hyper)graph complex, see for example the survey by Wachs [WAC03]. The majority of those results focuses on the graph case $d=2$ with varying $m$, and thus is somewhat complementary to our work (varying $d$, sometimes $m=2$). For $m=d=2$ all multigraded Betti numbers are known.
For $d=2$, the quadratic Veronese, Reiner and Roberts [RR00] studied the representation theoretic decomposition of the homology (in characteristic 0) of $\Delta_{\mathbf{b}}$. As a corollary, one obtains a formula for $\dim_{k}\tilde{H}_{p-1}(\Delta_{\mathbf{b}};k)$ in terms of Kostka numbers. By Hochster’s formula, this yields a formula for the multigraded Betti number $\beta_{p,\mathbf{b}}$ of the quadratic Veronese embedding.
Still for $d=2$, Dong [DON03] studied the homotopy type of bounded degree graph complexes and provided a complete characterization of when $\Delta_{\mathbf{b}}$ is contractible [DON03, Theorem 3.8]. For the special case of the quadratic Veronese, this contains our vanishing results. He also showed that under certain conditions $\Delta_{\mathbf{b}}$ is homotopy equivalent to a wedge of equidimensional spheres [DON03, Theorem 3.10]. The $d=2$ case of our Theorem 1.3 follows from Dong’s considerations.
For general $d$, Björner, Lovász, Vrećica and Živaljević [BLV+94] and Athanasiadis [ATH04, Theorem 1.2] studied the connectivity of the “hypergraph matching complex” $\Delta_{(1,\ldots,1)}$. They show that the complex is $\lfloor\frac{m-2d}{d+1}\rfloor$-connected. The weight space technique in [KRW01] can be used to lift this result to bounded degree hypergraph complexes: $\Delta_{\mathbf{b}}$ is $\lfloor\frac{|\mathbf{b}|-2d-1}{d+1}\rfloor$-acyclic over a field of characteristic 0. In the language of Betti numbers, this means that for each $1\leq q\leq m$, $\beta_{p,q}=0$ if $0\leq p\leq(q-1)d$. For the projective plane ($m=2$) this is a consequence of (1).
@@ Acknowledgements
We thank Milena Hering and the authors of [BEG+20] for enlightening conversations and encouragement. Both authors had the opportunity to discuss the present ideas during the Thematic Program in Commutative Algebra and Applications at the Fields Institute in Toronto, Canada. We also thank Vic Reiner for pointing us to bounded degree (hyper)graph complexes and the like.
@@ 2.1 Hochster’s formula
In this subsection, we provide background on the simplicial complex $\Delta_{\mathbf{b}}$ and Hochster’s formula. The references for these are [MS05, Section 9.1] and [STU96, Section 12.C].
@@
$\beta_{p,\mathbf{b}}=\dim_{k}\tilde{H}_{p-1}(\Delta_{\mathbf{b}};k)$.
@@ 2.2 Discrete Morse theory
In this subsection, we provide background on discrete Morse theory (cf. [FOR02, KOZ20]).
@@
A discrete Morse function on $\Delta$ [FOR02, Definition 2.1] is a function which, roughly speaking, assigns higher numbers to higher-dimensional simplices, with at most one exception, locally, at each simplex. Once a discrete Morse function $f$ is given, the discrete gradient vector field of $f$ is defined as a collection $W$ of pairs $\{\alpha^{(q)}<\beta^{(q+1)}\}$ of simplices, where $\{\alpha^{(q)}<\beta^{(q+1)}\}$ is in $W$ if and only if $f(\beta)\leq f(\alpha)$. A simplex is critical [FOR02, Definition 2.3] if and only if it is not in $W$.
Theorem 2.5 ([FOR02, Theorem 3.5]).
@@ 6 GENERALIZATION TO HIGHER DIMENSIONAL PROJECTIVE SPACE
$D\mathrel{\mathop{:}}=\Bigg\{I\in\binom{\{\binom{r+m-2}{m}+1,\dots,\binom{r+m-1}{m}\}}{p+1-\binom{r+m-2}{m}}\mid\hskip 216.81pt$
$\hskip 130.08731pt\mathbf{b}_{t}-\displaystyle\sum_{i=1}^{\binom{r+m-2}{m}}\mathbf{a}^{i}_{t}-1\leq\displaystyle\sum_{i\in I}\mathbf{a}^{i}_{t}\leq\mathbf{b}_{t}-\displaystyle\sum_{i=1}^{\binom{r+m-2}{m}}\mathbf{a}^{i}_{t}\quad\forall 1\leq t\leq m-1\Bigg\}$
and
@@ 7 Optimality of the Bounds $A_{p+1}$
The articles [EL12, EEL16] show that $\beta_{p,1}\neq 0$ for $1\leq p\leq\binom{d+m-1}{m}+m-2$. In the following, we give a combinatorial proof and show that, on the row $1$ of the Betti table, the bounds $A_{p+1}$ for the multigraded Betti numbers are sharp.